OpenStudy (anonymous):

Given log(base 6)8 is approx 1.1606 and log(base 7)9 is approx. 1.1292, what is the value of log(base 7)729

OpenStudy (anonymous):

log7(729) = log7(27^2) = 2*log7(27) = 2*log7(3^3) = 6*log7(3) since log7(9) = 2*log7(3), log7(3) = 1.1292/2 so, log7(729) = 6*1.1292/2